# Finite Distance Corrections to Vacuum Birefringence in Strong Gravitational and Electromagnetic Fields

### v1

@@ III.2 The finite-distance geometric angles
With the curvature of the optical manifolds established, we now define the geometric configuration of the lensing system at finite distances. We consider a light source located at a radial coordinate $r_{\text{S}}$ and a receiver (observer) located at $r_{\text{R}}$. In the non-asymptotically flat spacetime governed by the effective metrics derived in Section II, we cannot assume these points lie at infinity.
We define the local angle $\Psi^{\pm}$ as the angle between the tangent vector of the light ray, $K^{i}=dx^{i}/dt$, and the radial unit vector, $e_{rad}^{i}$, measured within the optical manifold $M^{opt}_{\pm}$. The local angle represents the deviation of the photon’s trajectory from the radial direction as measured by a local static observer. In terms of the optical metric components from Eq. (10), the cosine of this angle is given by the inner product $\cos\Psi^{\pm}=\gamma_{rr}^{\pm}(e_{rad}^{r}K^{r})$.
@@
$b=\frac{\gamma_{\phi\phi}^{\pm}}{\sqrt{\gamma_{ii}^{\pm}(dx^{i}/dt)(dx^{i}/dt)}}\frac{d\phi}{dt}\Big|_{\text{null}}=\frac{H_{\pm}(r)r^{2}}{A_{\pm}(r)}\frac{d\phi}{dt},$ (15)
where we have utilized the effective metric decomposition from Eq. (7).
@@
$\boxed{\sin\Psi_{\text{R,S}}^{\pm}=\frac{b}{r_{\text{R,S}}}\sqrt{\frac{A_{\pm}(r_{\text{R,S}})}{H_{\pm}(r_{\text{R,S}})}}.}$ (16)
Note that Equation (16) provides the magnitude of the sine of the angle. In the application of the Gauss-Bonnet theorem to the quadrilateral domain [Ishihara:2016vdc, Ishihara:2016sfv], $\Psi$ is defined as the angle between the outgoing radial unit vector and the photon tangent vector. For the receiver, this angle $\Psi_{\text{R}}$ is acute. However, for the source, the photon is directed inward toward the lens, making the angle obtuse. Therefore, in the sum $\Psi_{\text{R}}-\Psi_{\text{S}}$, the source angle is treated as $\Psi_{\text{S}}=\pi-\arcsin(\dots)$, consistent with the exterior angle formulation
Equation (16) highlights a critical consequence of the non-linear interaction: the geometric angles $\Psi_{R}^{\pm}$ and $\Psi_{S}^{\pm}$ are explicitly polarization-dependent. Unlike in standard General Relativity where the angle depends only on the gravitational potential $A(r)$, the NLED modification introduces the factor $H_{\pm}(r)$ in the denominator. Consequently, a receiver at a fixed coordinate $r_{\text{R}}$ observing a source at $r_{\text{S}}$ will measure two distinct image positions for the two polarization modes, even before calculating the integrated deflection along the path. Such a local birefringence at the endpoints contributes to the total observable splitting alongside the integrated curvature effect.
@@ III.3 The birefringence integral
$\iint_{{}^{\infty}_{R}\Box^{\infty}_{S}}K^{\pm}dS^{\pm}+\int_{\partial({}^{\infty}_{R}\Box^{\infty}_{S})}\kappa_{g}d\ell+\sum_{i}\theta_{i}=2\pi,$ (17)
@@
Following the geometric construction by Ishihara -et al. [Ishihara:2016vdc], the sum of the jump angles and the boundary integrals can be related to the geometric deflection angle $\alpha_{\pm}$. Specifically, the definition of the deflection angle $\alpha$ given in Eq. (1.17) of the reference text [Rindler:2007zz, Ishihara:2016vdc] (and generalized here for polarization modes) is:
$\alpha_{\pm}\equiv\Psi_{R}^{\pm}-\Psi_{S}^{\pm}+\phi_{RS}^{\pm},$ (18)
@@
$\boxed{\alpha_{\pm}=-\iint_{{}^{\infty}_{R}\Box^{\infty}_{S}}K^{\pm}dS^{\pm}.}$ (19)
@@
$\Delta\alpha=\alpha_{+}-\alpha_{-}=-\left(\iint_{D^{+}}K^{+}dS^{+}-\iint_{D^{-}}K^{-}dS^{-}\right),$ (20)
@@
Equation (20) presents a significant calculational challenge: the integration domains $D^{+}$ and $D^{-}$ are physically distinct because the photon paths $\gamma^{+}$ and $\gamma^{-}$ differ due to the effective metric splitting. To compute $\Delta\alpha$, one cannot simply subtract the integrands. Instead, we must account for both the difference in the curvature scalars ($K^{+}-K^{-}$) and the difference in the integration areas ($\delta S$). This requires a perturbative approach to map both domains onto a common reference manifold (the background geometry), which we will develop in the next section.
@@ IV Perturbative expansion and differential bending
In the previous section, we established that the finite-distance birefringence $\Delta\alpha$ is given by the difference of two area integrals over distinct domains in the optical manifolds $M^{opt}_{\pm}$. Direct integration of Eq. (20) is analytically intractable due to the complex dependencies of the effective metric functions $A_{\pm}$ and $H_{\pm}$ on the field invariants. To overcome this, we employ a perturbative approach. We assume the non-linear electrodynamics (NLED) coupling is weak, which is characterized by a small parameter $\beta$ (e.g., proportional to the fine-structure constant $\alpha_{\text{EM}}^{2}$ in QED), allowing us to linearize the geometric quantities around the background General Relativistic (GR) solution. The procedure decouples the gravitational deflection from the NLED-induced birefringence, enabling us to isolate the specific finite-distance corrections arising from the photon’s polarization.
@@ IV.1 The weak-field expansion
We begin by expanding the effective metric functions defined in Eq. (7) with respect to the NLED coupling parameter $\beta$. In the limit $\beta\to 0$, we recover the background metric functions $A(r)$ and $B(r)$, and the angular deformation factor $H(r)$ approaches unity. We write the expansions as:
$A_{\pm}(r)=A(r)\left(1+\beta\mathcal{A}_{\pm}(r)\right),\quad B_{\pm}(r)=B(r)\left(1+\beta\mathcal{B}_{\pm}(r)\right),\quad H_{\pm}(r)=1+\beta\mathcal{H}_{\pm}(r),$ (21)
where $\mathcal{A}_{\pm}$, $\mathcal{B}_{\pm}$, and $\mathcal{H}_{\pm}$ are dimensionless perturbation functions determined by the specific NLED Lagrangian (Eq. (5)).
Substituting these expansions into the general expression for the Gaussian curvature (Eq. (14)), we linearize $K^{\pm}(r)$. The curvature decomposes into the background gravitational curvature $K_{\text{GR}}(r)$ and a polarization-dependent correction:
$K^{\pm}(r)=K_{\text{GR}}(r)+\beta K^{(1)}_{\pm}(r)+O(\beta^{2}).$ (22)
Here, $K_{\text{GR}}$ is the Gaussian curvature of the optical manifold for the background metric (e.g., $K\approx-2M/r^{3}$ for Schwarzschild), and $K^{(1)}_{\pm}$ encapsulates the NLED contribution involving radial derivatives of $\mathcal{H}_{\pm}$.
To compute the integrals in Eq. (20), we must also define the integration domain boundaries. The upper boundary of the quadrilateral $^{\infty}_{R}\Box^{\infty}_{S}$ is defined by the photon’s spatial trajectory $r(\phi)$. Since the photons follow geodesics of different effective metrics, their trajectories differ. We parameterize the orbit using the inverse radial coordinate $u(\phi)\equiv 1/r(\phi)$. The orbit equation (corresponding to Eq. (11) in Ishihara -et al., modified here for the effective metric) is:
$\left(\frac{du}{d\phi}\right)^{2}+\frac{u^{2}H_{\pm}}{B_{\pm}}=\frac{1}{b^{2}}\frac{H_{\pm}}{A_{\pm}B_{\pm}}.$ (23)
@@
$u_{\pm}(\phi)=u_{0}(\phi)+\beta\delta u_{\pm}(\phi).$ (24)
Substituting the metric expansions (Eq. (21)) into the orbit equation (Eq. (23)) and keeping terms to first order in $\beta$, we find that the perturbation $\delta u_{\pm}$ satisfies the linearized Binet equation:
$\frac{d^{2}\delta u_{\pm}}{d\phi^{2}}+\delta u_{\pm}=S_{\pm}(\phi),$ (25)
@@
$\boxed{\delta u_{\pm}(\phi)=\int_{0}^{\phi}\sin(\phi-\varphi)\left[\frac{u_{0}(\varphi)}{2}\left(\mathcal{H}_{\pm}^{\prime}(u_{0})-\mathcal{A}_{\pm}^{\prime}(u_{0})-2\mathcal{B}_{\pm}(u_{0})\right)\right]d\varphi.}$ (26)
The integral in Eq. (26) describes the physical deviation of the photon path from the background null geodesic due to the effective NLED force. Crucially, this form remains regular at the turning point of the orbit, allowing for a consistent evaluation of the path-induced birefringence term.
The perturbation $\delta u_{\pm}$ is critical: it signifies that for a fixed coordinate angle $\phi$, the two polarization modes are located at different radial positions $r_{+}\neq r_{-}$. The spatial splitting implies that the integration domains $D^{+}$ and $D^{-}$ in Eq. (20) do not overlap perfectly. In the next subsection, we will utilize this trajectory shift to compute the area term of the differential bending angle.
@@ IV.2 The differential bending angle
With the trajectory perturbation $\delta u_{\pm}$ determined, we proceed to calculate the differential deflection angle $\Delta\alpha$, which constitutes the observable birefringence. As established in Eq. (20), $\Delta\alpha$ is defined by the difference between the Gauss-Bonnet integrals over the two polarization-dependent domains $D^{+}$ and $D^{-}$.
@@
$\Delta\alpha=-\left[\iint_{D_{0}}\left(K^{+}\sqrt{\gamma^{+}}-K^{-}\sqrt{\gamma^{-}}\right)drd\phi-\int_{\phi_{\text{S}}}^{\phi_{\text{R}}}\left(K_{\text{GR}}(r)\sqrt{\gamma_{\text{GR}}(r)}\right)\bigg|_{r=r_{0}(\phi)}\Delta r(\phi)d\phi\right],$ (27)
@@
The second term in Eq. (27) is the area shift correction. Using the perturbation relation $\Delta r\approx-\Delta(\delta u)/u_{0}^{2}$ derived from Eq. (26), we can express the total finite-distance birefringence explicitly in terms of the NLED perturbations. The formula isolates the NLED effects from the background gravitational lensing:
$\boxed{\Delta\alpha=-\underbrace{\int_{\phi_{\text{S}}}^{\phi_{\text{R}}}\int_{1/u_{0}(\phi)}^{\infty}\Delta\left(K\sqrt{\gamma}\right)drd\phi}_{\text{Refractive Curvature Difference}}-\underbrace{\int_{\phi_{\text{S}}}^{\phi_{\text{R}}}\left[\frac{K_{\text{GR}}}{u_{0}^{2}}\sqrt{\gamma_{\text{GR}}}\right]_{r=1/u_{0}}(\delta u_{+}-\delta u_{-})d\phi}_{\text{Path-Induced Birefringence}}.}$ (28)
The second term represents the geodesic shift contribution. It accounts for the fact that the two polarization modes traverse physically distinct regions of the background manifold. Because the background curvature $K_{\text{GR}}$ is spatially varying (non-uniform), the flux of curvature through the domain $D^{+}$ differs from that through $D^{-}$ simply because the boundaries $u_{+}(\phi)$ and $u_{-}(\phi)$ are not identical. Even if the effective NLED curvature correction $K^{(1)}$ were zero, this term would generate a non-zero birefringence $\Delta\alpha$ solely due to the splitting of the photon paths ($\delta u_{+}\neq\delta u_{-}$) within the strong gravitational gradient.
The above result is pivotal since it demonstrates that even if the local effective curvatures $K^{+}$ and $K^{-}$ were identical, a non-zero birefringence would still arise if the photon paths differ ($\delta u_{+}\neq\delta u_{-}$) within the background gravitational field. For finite distances, the limits of the angular integration $\phi_{\text{S}}$ and $\phi_{\text{R}}$ (corresponding to $r_{\text{S}}$ and $r_{\text{R}}$) are finite, meaning the birefringence accumulation is truncated relative to the asymptotic case. The truncation leads to the finite-distance corrections we aim to isolate in the final subsection.

### v2

@@ III.2 The finite-distance geometric angles
With the curvature of the optical manifolds established, we now define the geometric configuration of the lensing system at finite distances. We consider a light source located at a radial coordinate $r_{\text{S}}$ and a receiver (observer) located at $r_{\text{R}}$. Even when the background and effective geometries are asymptotically flat in the usual sense, realistic lensing/polarimetry setups place the source $S$ and receiver $R$ at finite radii $r_{S}$ and $r_{R}$, so the endpoint angles $\Psi_{S,R}$ must be evaluated at finite distance rather than by assuming $r_{S},r_{R}\to\infty$.
We define the local angle $\Psi^{\pm}$ as the angle between the tangent vector of the light ray, $K^{i}=dx^{i}/dt$, and the radial unit vector, $e_{rad}^{i}$, measured within the optical manifold $M^{opt}_{\pm}$. The local angle represents the deviation of the photon’s trajectory from the radial direction as measured by a local static observer. In terms of the optical metric components from Eq. (II.11), the cosine of this angle is given by the inner product $\cos\Psi^{\pm}=\gamma_{rr}^{\pm}(e_{rad}^{r}K^{r})$.
@@
$b=\frac{\gamma_{\phi\phi}^{\pm}}{\sqrt{\gamma_{ii}^{\pm}(dx^{i}/dt)(dx^{i}/dt)}}\frac{d\phi}{dt}\Big|_{\text{null}}=\frac{H_{\pm}(r)r^{2}}{A_{\pm}(r)}\frac{d\phi}{dt},$ (III.5)
where we have utilized the effective metric decomposition from Eq. (II.8).
@@
$\sin\Psi_{\text{R,S}}^{\pm}=\frac{b}{r_{\text{R,S}}}\sqrt{\frac{A_{\pm}(r_{\text{R,S}})}{H_{\pm}(r_{\text{R,S}})}}.$ (III.6)
Note that Equation (III.6) provides the magnitude of the sine of the angle. In the application of the Gauss-Bonnet theorem to the quadrilateral domain [45, 44], $\Psi$ is defined as the angle between the outgoing radial unit vector and the photon tangent vector. For the receiver, this angle $\Psi_{\text{R}}$ is acute. However, for the source, the photon is directed inward toward the lens, making the angle obtuse. Therefore, in the sum $\Psi_{\text{R}}-\Psi_{\text{S}}$, the source angle is treated as $\Psi_{\text{S}}=\pi-\arcsin(\dots)$, consistent with the exterior angle formulation
Equation (III.6) highlights a critical consequence of the non-linear interaction: the geometric angles $\Psi_{R}^{\pm}$ and $\Psi_{S}^{\pm}$ are explicitly polarization-dependent. Unlike in standard General Relativity where the angle depends only on the gravitational potential $A(r)$, the NLED modification introduces the factor $H_{\pm}(r)$ in the denominator. Consequently, a receiver at a fixed coordinate $r_{\text{R}}$ observing a source at $r_{\text{S}}$ will measure two distinct image positions for the two polarization modes, even before calculating the integrated deflection along the path. Such a local birefringence at the endpoints contributes to the total observable splitting alongside the integrated curvature effect.
@@ III.3 The birefringence integral
$\iint_{{}^{\infty}_{R}\Box^{\infty}_{S}}K^{\pm}dS^{\pm}+\int_{\partial({}^{\infty}_{R}\Box^{\infty}_{S})}\kappa_{g}d\ell+\sum_{i}\theta_{i}=2\pi,$ (III.7)
@@
Following the geometric construction by Ishihara et al. [45], the sum of the jump angles and the boundary integrals can be related to the geometric deflection angle $\alpha_{\pm}$. Specifically, the definition of the deflection angle $\alpha$ given in Eq. (1.17) of the reference text [75, 45, 83] (and generalized here for polarization modes) is:
$\alpha_{\pm}\equiv\Psi_{R}^{\pm}-\Psi_{S}^{\pm}+\phi_{RS}^{\pm},$ (III.8)
@@
$\alpha_{\pm}=-\iint_{{}^{\infty}_{R}\Box^{\infty}_{S}}K^{\pm}dS^{\pm}.$ (III.9)
@@
$\Delta\alpha=\alpha_{+}-\alpha_{-}=-\left(\iint_{D^{+}}K^{+}dS^{+}-\iint_{D^{-}}K^{-}dS^{-}\right),$ (III.10)
@@
Equation (III.10) presents a significant calculational challenge: the integration domains $D^{+}$ and $D^{-}$ are physically distinct because the photon paths $\gamma^{+}$ and $\gamma^{-}$ differ due to the effective metric splitting. To compute $\Delta\alpha$, one cannot simply subtract the integrands. Instead, we must account for both the difference in the curvature scalars ($K^{+}-K^{-}$) and the difference in the integration areas ($\delta S$). This requires a perturbative approach to map both domains onto a common reference manifold (the background geometry), which we will develop in the next section.
@@ IV Perturbative expansion and differential bending
In the previous section, we established that the finite-distance birefringence $\Delta\alpha$ is given by the difference of two area integrals over distinct domains in the optical manifolds $M^{opt}_{\pm}$. Direct integration of Eq. (III.10) is analytically intractable due to the complex dependencies of the effective metric functions $A_{\pm}$ and $H_{\pm}$ on the field invariants. To overcome this, we employ a perturbative approach. We assume the NLED coupling is weak, which is characterized by a small parameter $\beta$ (e.g., proportional to the fine-structure constant $\alpha_{\text{EM}}^{2}$ in QED), allowing us to linearize the geometric quantities around the background General Relativistic (GR) solution. The procedure decouples the gravitational deflection from the NLED-induced birefringence, enabling us to isolate the specific finite-distance corrections arising from the photon’s polarization.
@@ IV.1 The weak-field expansion
We begin by expanding the effective metric functions defined in Eq. (II.8) with respect to the NLED coupling parameter $\beta$. In the limit $\beta\to 0$, we recover the background metric functions $A(r)$ and $B(r)$, and the angular deformation factor $H(r)$ approaches unity. We write the expansions as:
$A_{\pm}(r)=A(r)\Big(1+\beta\,a_{\pm}(r)\Big),\qquad B_{\pm}(r)=B(r)\Big(1+\beta\,b_{\pm}(r)\Big),\qquad H_{\pm}(r)=1+\beta\,h_{\pm}(r),$ (IV.1)
where $a_{\pm}$, $b_{\pm}$, and $h_{\pm}$ are dimensionless $\mathcal{O}(1)$ perturbation functions determined by the specific NLED Lagrangian (Eq. (II.5)). (The symbol $b_{\pm}$ should not be confused with the impact parameter $b$.)
Substituting these expansions into the general expression for the Gaussian curvature (Eq. (III.4)), we linearize $K^{\pm}(r)$. The curvature decomposes into the background gravitational curvature $K_{\text{GR}}(r)$ and a polarization-dependent correction:
$K^{\pm}(r)=K_{\text{GR}}(r)+\beta K^{(1)}_{\pm}(r)+O(\beta^{2}).$ (IV.2)
Here, $K_{\text{GR}}$ is the Gaussian curvature of the optical manifold for the background metric (e.g., $K\approx-2M/r^{3}$ for Schwarzschild), and $K^{(1)}_{\pm}$ encapsulates the NLED contribution involving radial derivatives of the perturbation functions $a_{\pm}$, $b_{\pm}$, and $h_{\pm}$.
To compute the integrals in Eq. (III.10), we must also define the integration domain boundaries. The upper boundary of the quadrilateral $^{\infty}_{R}\Box^{\infty}_{S}$ is defined by the photon’s spatial trajectory $r(\phi)$. Since the photons follow geodesics of different effective metrics, their trajectories differ. We parameterize the orbit using the inverse radial coordinate $u(\phi)\equiv 1/r(\phi)$. The orbit equation (corresponding to Eq. (11) in Ishihara et al., modified here for the effective metric) is:
$\left(\frac{du}{d\phi}\right)^{2}+\frac{u^{2}H_{\pm}}{B_{\pm}}=\frac{1}{b^{2}}\frac{H_{\pm}}{A_{\pm}B_{\pm}}.$ (IV.3)
@@
$u_{\pm}(\phi)=u_{0}(\phi)+\beta\delta u_{\pm}(\phi).$ (IV.4)
Substituting the metric expansions (Eq. (IV.1)) into the orbit equation (Eq. (IV.3)) and keeping terms to first order in $\beta$, we find that the perturbation $\delta u_{\pm}$ satisfies the linearized Binet equation:
$\frac{d^{2}\delta u_{\pm}}{d\phi^{2}}+\delta u_{\pm}=S_{\pm}(\phi),$ (IV.5)
@@
$\delta u_{\pm}(\varphi)=\int_{0}^{\varphi}\sin(\varphi-\phi)\,\Big[u_{0}(\phi)^{2}\big(h_{\pm}^{\prime}(u_{0})-a_{\pm}^{\prime}(u_{0})-2\,b_{\pm}(u_{0})\big)\Big]\;d\phi.$ (IV.6)
Here a prime denotes differentiation with respect to $u$ evaluated along the background orbit $u=u_{0}(\phi)$. The integral in Eq. (IV.6) describes the physical deviation of the photon path from the background null geodesic due to the effective NLED force. Crucially, this form remains regular at the turning point of the orbit, allowing for a consistent evaluation of the path-induced birefringence term.
The perturbation $\delta u_{\pm}$ is critical: it signifies that for a fixed coordinate angle $\phi$, the two polarization modes are located at different radial positions $r_{+}\neq r_{-}$. The spatial splitting implies that the integration domains $D^{+}$ and $D^{-}$ in Eq. (III.10) do not overlap perfectly. In the next subsection, we will utilize this trajectory shift to compute the area term of the differential bending angle.
@@ IV.2 The differential bending angle
With the trajectory perturbation $\delta u_{\pm}$ determined, we proceed to calculate the differential deflection angle $\Delta\alpha$, which constitutes the observable birefringence. As established in Eq. (III.10), $\Delta\alpha$ is defined by the difference between the Gauss-Bonnet integrals over the two polarization-dependent domains $D^{+}$ and $D^{-}$.
@@
$\Delta\alpha=-\left[\iint_{D_{0}}\left(K^{+}\sqrt{\gamma^{+}}-K^{-}\sqrt{\gamma^{-}}\right)drd\phi-\int_{\phi_{\text{S}}}^{\phi_{\text{R}}}\left(K_{\text{GR}}(r)\sqrt{\gamma_{\text{GR}}(r)}\right)\bigg|_{r=r_{0}(\phi)}\Delta r(\phi)d\phi\right],$ (IV.7)
@@
The second term in Eq. (IV.7) is the area shift correction. Using the perturbation relation $\Delta r\approx-\Delta(\delta u)/u_{0}^{2}$ derived from Eq. (IV.6), we can express the total finite-distance birefringence explicitly in terms of the NLED perturbations. The formula isolates the NLED effects from the background gravitational lensing:
$\Delta\alpha=-\underbrace{\int_{\phi_{\text{S}}}^{\phi_{\text{R}}}\int_{1/u_{0}(\phi)}^{\infty}\Delta\left(K\sqrt{\gamma}\right)drd\phi}_{\text{Refractive Curvature Difference}}-\underbrace{\int_{\phi_{\text{S}}}^{\phi_{\text{R}}}\left[\frac{K_{\text{GR}}}{u_{0}^{2}}\sqrt{\gamma_{\text{GR}}}\right]_{r=1/u_{0}}(\delta u_{+}-\delta u_{-})d\phi}_{\text{Path-Induced Birefringence}}.$ (IV.8)
The second term in Eq. (IV.8) is a boundary (domain-shift) correction: it arises when the two Gauss-Bonnet area integrals over $D_{+}$ and $D_{-}$ are mapped onto the same reference domain $D_{0}$ bounded by the background orbit $r_{0}(\phi)=1/u_{0}(\phi)$. Concretely, for any smooth integrand $\mathcal{I}(r,\phi)$ one has, by the Leibniz rule,
$\int_{\phi_{S}}^{\phi_{R}}\!\!\int_{r_{\pm}(\phi)}^{\infty}\mathcal{I}(r,\phi)\,dr\,d\phi=\int_{\phi_{S}}^{\phi_{R}}\!\!\int_{r_{0}(\phi)}^{\infty}\mathcal{I}(r,\phi)\,dr\,d\phi-\int_{\phi_{S}}^{\phi_{R}}\left.\mathcal{I}(r,\phi)\right|_{r=r_{0}(\phi)}\delta r_{\pm}(\phi)\,d\phi+\mathcal{O}(\beta^{2}),$ (IV.9)
with $\delta r_{\pm}(\phi)\equiv r_{\pm}(\phi)-r_{0}(\phi)\simeq-\delta u_{\pm}(\phi)/u_{0}(\phi)^{2}$. Applying Eq. (IV.9) to $\mathcal{I}=K\sqrt{\gamma}$ and subtracting the $\pm$ expressions reproduces Eq. (IV.8) identically to $\mathcal{O}(\beta)$. In this sense, the split in Eq. (IV.8) is a convenient bookkeeping device tied to the choice of reference curve $u_{0}$: each term separately changes if $u_{0}$ is shifted by a common $\mathcal{O}(\beta)$ amount, but their sum (the observable $\Delta\alpha$) is invariant to the working order.
Physically, the boundary term measures the change in the background-curvature flux induced by the polarization-dependent displacement of the photon orbit. A nonzero contribution therefore requires $\delta u_{+}\neq\delta u_{-}$, i.e. genuinely distinct polarization-dependent geodesics. While one can construct special cases in which the bulk integrand difference $\Delta(K\sqrt{\gamma})$ is suppressed on $D_{0}$ yet $\delta u_{+}-\delta u_{-}\neq 0$, this requires a non-generic cancellation in the curvature/area densities; in typical NLED models the same mode dependence that produces trajectory splitting also yields $\Delta(K\sqrt{\gamma})\neq 0$. In particular, for an exceptional non-birefringent theory in which $g^{\mu\nu}_{\rm eff,+}$ and $g^{\mu\nu}_{\rm eff,-}$ coincide (up to a common conformal rescaling) one has $\delta u_{+}=\delta u_{-}$ and hence the boundary contribution vanishes identically; Born-Infeld electrodynamics is of this type [13, 67]. For finite distances, the limits of the angular integration $\phi_{\text{S}}$ and $\phi_{\text{R}}$ (corresponding to $r_{\text{S}}$ and $r_{\text{R}}$) are finite, meaning the birefringence accumulation is truncated relative to the asymptotic case. The truncation leads to the finite-distance corrections we aim to isolate in the final subsection.